During March 2019, Alaska Corporation recorded $269,000 of costs related to factory overhead.  Alaska's overhead application rate is based on direct labor hours. The preset formula for overhead application estimated that $250,000 would be incurred, and 12,500 direct labor hours would be worked.  During March, 12,000 hours were actually worked. Use this information to determine the amount of overhead over or under applied.  Enter overapplied overhead as a negative number. (round & enter any final dollar answers to the nearest whole dollar)
Answer #1

Predetermined Overhead rate = Estimated overhead/Estimated total labor hours = 250,000/12,500 = 20 per dlh

Applied overhead = 20*12,000 = 240,000

Actual overhead = 269,000

So, overhead is underapplied by 29,000